OpenStudy (highschoolmom2010):

What is the graph of the function? f(x)=1/3x^2

OpenStudy (anonymous):

a.to find the inverse of f(x) we shall replace the value of x by y and vise versa. then the inverse of y=1/3x+2 will be x=1/3y+2 and when we solve for y we will find that y=3x-6 and when we replace y by g(x) the inverse will be g(x)=3x-6 b. f(g(x)=g(f(x) then 1/3(3x-6)+2=3(1/3x+2)-6 when we solve this we found that x=x. c. the graph for f(x) is an oblique line that passes through the points(-6,2) and the graph for g(x) is an oblique line which passes through the points(2,-6)

OpenStudy (highschoolmom2010):

what.....

OpenStudy (anonymous):

this shold help you on this

OpenStudy (highschoolmom2010):

those are my choices because i have tried doing this on my own and i dont get it at all

OpenStudy (aum):

The question is what is the graph of 1/3 * x^2 not to find the inverse.

OpenStudy (aum):

f(x) = 1/2 * x^2 This is the equation of a vertical parabola. Since the coefficient of x^2 is positive, it is a parabola that opens upward. The vertex of this parabola is the origin. With these three information you can identify the parabola in the answer choices.

OpenStudy (aum):

Which parabola has the vertex at the origin?

OpenStudy (highschoolmom2010):

the second one

OpenStudy (aum):

Correct. That is the answer.

OpenStudy (anonymous):

how

OpenStudy (highschoolmom2010):

but it wants 1/3* x^2 thats where im confused now

OpenStudy (aum):

Yeah, that was my typo but the explanation is the same whether it is 1/2 * x^2 or 1/3 * x^2. For both the vertex is the origin. Among the answer choices there is only one that has the origin as the vertex.

OpenStudy (highschoolmom2010):

oh ok

OpenStudy (aum):

In a vertical parabola, the vertex is the lowermost point or in the case of a parabola that opens downward, the vertex is the uppermost point. In this case the parabola opens upward and the vertex is the lowermost point. Only the second one has vertex at the origin.

OpenStudy (highschoolmom2010):

thanks for explaining

OpenStudy (aum):

You are welcome.